OpenStudy (anonymous):

Write the equation of the line in slope-intercept form with a slope of 2 and a y-intercept of -3

OpenStudy (calculusfunctions):

Equation of any diagonal (oblique) line is y=mx+b where m is the slope and b is the y-intercept.

OpenStudy (ajprincess):

Slope-intercept form is y=mx+c m-slope c-y-intercept To find the eqn put 2 in place of m and -3 in place of y in y=mx+c

OpenStudy (anonymous):

y=2x-3 y=mx+b

OpenStudy (calculusfunctions):

Yes!!

OpenStudy (anonymous):

y = mx + b (slope intercept form) y = 2x - 3 :)